# Mass and energy cascade in collisionless dark matter flow and relevant constraints on the dark matter particle massA.6

### v3

@@ A.2 Time scales in mass cascade
which is the average waiting time (lifespan) for a given halo in a halo group of mass scale $m_{h}$, or equivalently the time required to cascade the entire mass $m_{g}$ of that halo group. The time scale $\tau_{g}$ should decrease with $m_{h}$ due to faster mass accretion of larger haloes.
@@
We are now ready to determine the characteristic mass scale $m_{h}^{*}$ for mass cascade. Let $\tau_{g}\left(m_{h},a\right)$ be the average waiting time for a halo of mass $m_{h}$ to merge with a single merger of mass $m_{p}$ at physical time t. The fourth time scale $\tau_{f}\left(m_{h},a\right)$ that we will introduce is
@@
For small haloes with mass $m_{h}<m_{h}^{*}$, we expect the time scale $\tau_{f}\left(m_{h},a\right)\ll t$ to have sufficient time to form these haloes. For haloes with mass $m_{h}>m_{h}^{*}$, the time required to form that halo $\tau_{f}\left(m_{h},a\right)\gg t$ and these haloes are very rare to find at current time t. The time required to form haloes of a characteristic mass $m_{h}^{*}$ should be exactly on the order of the current physical time t, i.e.$\tau_{f}\left(m_{h}^{*},a\right)\sim t\sim\tau_{M}\left(a\right)$, from which we can derive,
$-\frac{m_{h}^{*}}{\varepsilon_{m}}n_{h}^{*}n_{p}^{*}\sim-\frac{M_{h}\left(a\right)}{\varepsilon_{m}}\sim\frac{1}{H},$ (119)
@@
In short, two distinct ranges can be identified for inverse mass cascade from time/mass scales: 1) mass propagation range with $m_{h}<m_{h}^{*}$, where the system is in a statistically steady state with a scale-independent mass flux $\varepsilon_{m}$ and a time-invariant group mass $m_{g}=m_{h}n_{h}$ (Fig. 20); 2) mass deposition range with $m_{h}>m_{h}^{*}$, where mass cascaded from small scales is actively consumed to grow haloes. Halo group mass $m_{g}$ is increasing with time in this range.
A.3 Mass flux and mass transfer functions
@@
[…]
Here $M_{h}\left(a\right)$ is the total mass in the halo sub-system that increases with scale factor a. The halo mass function $f_{M}\left(m_{h},a\right)$ is the probability distribution of total mass $M_{h}\left(a\right)$ with respect to halo mass $m_{h}$.
Since halo mass $m_{h}$ and scale factor a are the only two independent variables, the mass function can be written as a function of $m_{h}$ and a, i.e. $f_{M}\left(m_{h},a\right)=f_{M}\left(m_{h},m_{h}^{*}\left(a\right)\right)$. The characteristic mass scale $m_{h}^{*}\left(a\right)$ varies with the scale factor a only, a monotonically increasing function reflecting the fact that larger haloes emerge at a later time. The mass flux function $\Pi_{m}$ across halo groups should be independent of halo size $m_{h}$ for halo groups smaller than $m_{h}^{*}\left(a\right)$, where the mass flux function reduces to
@@
The constant mass flux (or the mass dissipation rate $\varepsilon_{m}$ that is independent of mass scale $m_{h}$) cascades mass from the smallest mass scale to the characteristic scale ($0\ll m_{h}<m_{h}^{*}$) in the mass propagation range. The total mass of the halo group of size $m_{h}$ is
@@
Figure 20: The variation of halo group mass $m_{g}$ (normalized by particle mass $m_{p}$) with group size $n_{p}$ at different redshift $z$. At a statistically steady state, $m_{g}$ does not vary with time for halo groups with mass below $m_{h}*$, where mass propagation is dominant. Data for $m_{g}$ is from halo-based statistical dataset for dark matter flow (Xu, 2022c) and is also used to compute rate of mass cascade $\Pi_{m}$ in Fig. 4.
The real-space mass transfer function can be defined as the derivative of the mass flux function with respect to halo mass,
@@
which quantifies the rate of change of group mass $m_{g}\left(m_{h},a\right)$. For the mass propagation range,
@@
The mass transfer function $T_{m}\left(m_{h},a\right)$ describes the removal of mass from a small scale and the deposition of mass at a large scale ($T_{m}\left(m_{h},a\right)>0$ for $m_{h}>m_{h}^{*}$).
@@
Let the time scale $\tau_{h}\left(m_{h},a\right)$ be the average time for a single merging event in a halo group of mass $m_{h}$, or equivalently an event frequency $f_{h}\left(m_{h},a\right)$. The rate of mass transfer from the scale below $m_{h}$ to the scale above $m_{h}$ is,
@@
where $\alpha_{0}$ is a numerical factor on the order of unity. The event frequency $f_{h}\left(m_{h},a\right)$ should be proportional to the number of haloes in the group (term 1 in Eq. (129)) and the surface area of the haloes (term 2). Because the halo interactions in mass space are local, we can assume the mass cascade involves merging events between a halo of similar size and a single merger (Figs. 17 and 18). Halo group with more haloes (term 1 in Eq. (129)) and haloes with a larger surface area (proportional to $m_{h}^{\lambda}$, i.e., the term 2 in Eq. (129)) have a greater probability of merging with a single merger,
@@
where $f_{0}\left(a\right)$ is a fundamental frequency for the merging between two single mergers at a given redshift z or scale factor a and may be used to determine dark matter particle mass $m_{p}$ (Eq. (139)). The characteristic time of a single merging event can be written as $\tau_{h}\left(m_{h},a\right)={1/\left(\alpha_{0}f_{h}\right)}$.
Without loss of generality, the exponent $\lambda$ is a halo geometry parameter that represents the effect of halo surface area on the merging frequency $f_{h}\left(m_{h},a\right)$. For two haloes of very different sizes (merging between a large halo and a single merger), it is estimated that $\lambda={2/3}$ with $m_{h}\propto r_{h}^{3}\propto A_{h}^{{3/2}}$, where $A_{h}$ is the halo surface area. For small haloes, merging is more likely between two haloes of comparable sizes where $\lambda$ can deviate from 2/3 and approach 1.
Substitution of Eq. (129) for the event frequency into the Eq. (128) leads to the mass flux,
@@
For self-similar gravitational clustering in the mass propagation range ($m_{h}<m_{h}^{*}$), the halo mass $m_{h}$ and characteristic mass scale $m_{h}^{*}$ are the only two controlling variables. We can simply express the mass function as $f_{M}\left(m_{h},m_{h}^{*}\right)\sim\left(m_{h}\right)^{x}\left(m_{h}^{*}\right)^{-x-1}$. Now using dimensional analysis, the only possible form of the mass function $f_{M}$ that satisfies Eq. (130) is ($f_{M}$ should have a unit of 1/kg in SI units and $\varepsilon_{m}$ is a function of a only and is independent of $m_{h}$),
@@
which is a time-invariant constant and dependent only on the mass resolution $m_{p}$. For $m_{p}=2.27\times 10^{11}{M_{\odot}/h}$ from Table 1 and $m_{h}^{*}\left(z=0\right)\approx 2\times 10^{13}{M_{\odot}/h}$, $\lambda_{0}$ is on the order of 0.13.
@@
that is also related to the Hubble constant $H\left(a\right)$ and mass resolution $m_{p}$. The characteristic time scale $\tau_{h}^{*}=\tau_{h}\left(m_{h}^{*},a\right)$ associated with the characteristic mass $m_{h}^{*}$ is
@@
The fundamental frequency $f_{0}\left(a\right)$ is the frequency for the elementary merging between two single mergers and is expected to decrease with time. Without loss of generality, let’s assume a power-law of $f_{0}\left(a\right)\propto a^{-\tau_{0}}$ that leads to $\varepsilon_{m}\left(a\right)\propto a^{-\tau_{0}}$ (Eq. (132)), $\left(m_{h}^{*}\right)^{1-\lambda}\propto a^{{3/2}-\tau_{0}}$ (Eq. (138)), and $M_{h}\left(a\right)\propto a^{{3/2}-\tau_{0}}$ (Eq. (134)). Once the statistically steady state is established for inverse mass cascade, the total mass of the halo sub-system increases as $M_{h}\left(a\right)\propto a^{{3/2}-\tau_{0}}$ regardless of the value of $\lambda$. Obviously,
@@
With total mass in halo sub-system $M_{h}\left(a\right)$ increasing with the scale factor a, it requires $0<\tau_{0}<{3/2}$. With $\tau_{0}>0$ and $a\to\infty$, the mass flux $\varepsilon_{m}\left(a\right)$ approaches zero with the entire system approaches the limiting thermodynamic equilibrium but can never reach. In this regard, the inverse mass cascade with a constant rate is a key feature of the statistically (intermediate) steady state when the system evolves toward the limiting equilibrium.
Number of haloes $n_{h}\left(m_{h}\right)$ in halo group with a given mass $m_{h}$ is
@@
, which does not vary with time once the statistically steady state is established. Substitution of Eq. (141) into Eq. (139), we can express the fundamental frequency $f_{0}\left(a\right)$ as,
@@
which is dependent on the mass resolution $m_{p}$ only (the mass of dark matter particle), with smaller particle mass $m_{p}$ giving rise to a greater fundamental frequency $f_{0}\left(a\right)$ in N-body simulation. For mass resolution of $m_{p}=2.27\times 10^{11}{M_{\odot}/h}$ from Table 1 and $m_{h}^{*}\left(a=1\right)=2\times 10^{13}{M_{\odot}/h}$, $b_{0}\approx 2.2$ with $\tau_{0}=1$. In other words, the dark matter particle mass can be determined if the fundamental frequency $f_{0}\left(a\right)$ can be precisely measured.
@@
which is the mass fraction of halo sub-system $M_{h}\left(a\right)$ and not dependent on the mass resolution $m_{p}$ and scale factor a. We estimate $c_{0}\approx 0.29$ with $\tau_{0}=1$, i.e. $M_{h}\approx 0.58Nm_{p}$ when a=1. This information is used to study the density distributions for particles in haloes and out-of-haloes, respectively (Xu, 2022g).
Finally we present the simplified expressions for time and mass scales and mass flux function that can be fully described as functions of halo mass $m_{h}$, scale factor a, and mass resolution $m_{p}$ with four numerical constants $\beta_{0}$ (pre-factor for halo mass function $f_{M}$), $c_{0}$ (the mass fraction of halo sub-system), $\lambda$ (the halo geometry parameter), and $\tau_{0}$ (the exponent of the fundamental frequency $f_{0}$), and the characteristic mass $m_{h}^{*}\left(a=1\right)$ and $H_{0}$,
@@
Let us revisit the time scales we introduced in Section A.2. The time scale $\tau_{M}\left(a\right)$ is
@@
Time scale $\tau_{h}$ for a single merging in halo group of mass $m_{h}$ is,
@@
The time scale $\tau_{g}$ that takes to cascade the group mass $m_{g}$ for the halo group of mass $m_{h}$ is,
@@
As expected, the mean waiting time (lifespan) $\tau_{g}$ of a given halo decreases with halo mass as $\tau_{g}\propto m_{h}^{-\lambda}$ and increases with a. Larger haloes have a shorter lifespan. Extremely large haloes have very fast mass accretion and an infinitesimal lifespan.
@@
which increases with halo mass as $\tau_{f}\propto m_{h}^{1-\lambda}a^{\tau_{0}}$ such that Smaller haloes were formed earlier with smaller $\tau_{f}$. The relation between time scales $\tau_{f}\left(m_{h}^{*},a\right)=\beta_{0}\tau_{M}\left(a\right)$ can be easily obtained from Eqs. (150) and (153) that is consistent with our analysis in Section A.2 (Eq. (120)), i.e. $\beta_{0}M_{h}\left(a\right)=m_{g}^{*}n_{p}^{*}=n_{h}^{*}n_{p}^{*2}m_{p}$ and $z_{h}^{*}\beta_{0}=1$ (from Eq. (121)). Different time scales can be related to the mass flux as
@@
The corresponding time scales at characteristic mass $m_{h}^{*}$ are
@@
The two time scales $\tau_{g}\propto a^{\tau_{0}}m_{h}^{-\lambda}$ (Eq. (152)) and $\tau_{f}\propto a^{\tau_{0}}m_{h}^{1-\lambda}$ (Eq. (153)), where larger haloes have a shorter lifespan but take more time to form. For $\lambda={2/3}$ and $\tau_{0}=1$, the lifespan $\tau_{g}^{*}$ of characteristic haloes is independent of time that is consistent with Eq. (24).
@@
which follows the same scaling as characteristic mass scale $m_{h}^{*}\left(a\right)$. For large haloes with $\tau_{0}=1$ and $\lambda={2/3}$, $m_{h}^{L}\left(a\right)\sim m_{h}^{*}\left(a\right)\sim a^{{3/2}}$. This scaling matches the mass growth of type II haloes, i.e., the dominant type for large haloes (see McBride et al., 2009, Fig. 2), as shown in Fig. 21.
Figure 21: The halo mass (normalized by $10^{12}M_{\odot}$) accretion history for type II haloes, i.e. the dominant type for large haloes (see McBride et al., 2009, Fig. 2), exhibits a power law scaling $\propto a^{{3/2}}$.
With the help of Eq. (117), we confirm that time scale $\tau_{f}$ (Eq. (153)) to form the typical halo is on the same order of t (Eq. (117)),
@@
A.4 The probability distribution of waiting time

### v4

@@ C.2 Time scales in mass cascade
which is the average waiting time (lifespan) for a given halo in a halo group of mass scale $m_{h}$, or equivalently, the time required to cascade the entire mass $m_{g}$ of that halo group. The time scale $\tau_{g}$ should decrease with $m_{h}$ due to faster mass accretion of larger haloes.
@@
We are now ready to determine the characteristic mass scale $m_{h}^{*}$ for the mass cascade. Let $\tau_{g}\left(m_{h},a\right)$ be the average waiting time for a halo of mass $m_{h}$ to merge with a single merger of mass $m_{p}$ at physical time t. The fourth time scale $\tau_{f}\left(m_{h},a\right)$ that we will introduce is
@@
[…]
$\frac{m_{h}^{*}}{\varepsilon_{m}}n_{h}^{*}n_{p}^{*}\sim-\frac{M_{h}\left(a\right)}{\varepsilon_{m}}\sim\frac{1}{H},$ (135)
@@
In short, two distinct ranges can be identified for inverse mass cascade from time/mass scales: 1) mass propagation range with $m_{h}<m_{h}^{*}$, where the system is in a statistically steady state with a scale-independent mass flux $\varepsilon_{m}$ and a time-invariant group mass $m_{g}=m_{h}n_{h}$ (Fig. 22); 2) mass deposition range with $m_{h}>m_{h}^{*}$, where mass cascaded from small scales is actively consumed to grow haloes. The mass of the halo group $m_{g}$ increases with time in this range.
C.3 Mass flux and mass transfer functions
@@
$\Pi_{m}\left(m_{h},a\right)=-\frac{\partial}{\partial t}\left[M_{h}\left(a\right)\int_{m_{h}}^{\infty}f_{M}\left(m,m_{h}^{*}\right)dm\right].$ (138)
Here $M_{h}\left(a\right)$ is the total mass in the halo sub-system that increases with scale factor a. The halo mass function $f_{M}\left(m_{h},a\right)$ is the probability distribution of the total mass $M_{h}\left(a\right)$ with respect to the halo mass $m_{h}$.
Since the halo mass $m_{h}$ and the scale factor a are the only two independent variables, the mass function can be written as a function of $m_{h}$ and a, i.e., $f_{M}\left(m_{h},a\right)=f_{M}\left(m_{h},m_{h}^{*}\left(a\right)\right)$. The characteristic mass scale $m_{h}^{*}\left(a\right)$ varies with the scale factor a only, a monotonically increasing function reflecting the fact that larger haloes emerge at a later time. The mass flux function $\Pi_{m}$ across the halo groups should be independent of halo size $m_{h}$ for halo groups smaller than $m_{h}^{*}\left(a\right)$, where the mass flux function reduces to
@@
The constant mass flux (or the mass dissipation rate $\varepsilon_{m}$ that is independent of mass scale $m_{h}$) cascades mass from the smallest mass scale to the characteristic scale ($0\ll m_{h}<m_{h}^{*}$) in the mass propagation range. The total mass of the halo group $m_{h}$ is
@@
Figure 22: The variation of halo group mass $m_{g}$ (normalized by particle mass $m_{p}$) with group size $n_{p}$ at different redshift $z$. At a statistically steady state, $m_{g}$ does not vary with time for halo groups with mass below $m_{h}*$, where mass propagation is dominant. Data for $m_{g}$ is from halo-based statistical dataset for dark matter flow (Xu:2022-Dark_matter-flow-dataset-part1) and is also used to compute rate of mass cascade $\Pi_{m}$ in Fig. 8.
The real-space mass transfer function can be defined as the derivative of the mass flux function with respect to the halo mass;
@@
which quantifies the rate of change in the mass of the group $m_{g}\left(m_{h},a\right)$. For the mass propagation range,
@@
The mass transfer function $T_{m}\left(m_{h},a\right)$ describes the removal of mass on a small scale and the deposition of mass on a large scale ($T_{m}\left(m_{h},a\right)>0$ for $m_{h}>m_{h}^{*}$).
@@
Let the time scale $\tau_{h}\left(m_{h},a\right)$ be the average time for a single merging event in a halo group of mass $m_{h}$, or equivalently an event frequency $f_{h}\left(m_{h},a\right)$. The rate of mass transfer from the scale below $m_{h}$ to the scale above $m_{h}$ is
@@
where $\alpha_{0}$ is a numerical factor on the order of unity. The frequency of the event $f_{h}\left(m_{h},a\right)$ should be proportional to the number of haloes in the group (term 1 in Eq. (145)) and the surface area of the haloes (term 2). Because the halo interactions in mass space are local, we can assume that the mass cascade involves merging events between a halo of similar size and a single merger (Figs. 19 and 20). The halo group with more haloes (term 1 in Eq. (145)) and haloes with a larger surface area (proportional to $m_{h}^{\lambda}$, i.e., term 2 in Eq. (145)) have a greater probability of merging with a single merger,
@@
where $f_{0}\left(a\right)$ is a fundamental frequency for the merging between two single mergers at a given redshift z or scale factor a and may be used to determine the mass of dark matter particles $m_{p}$ (Eq. (155)). The characteristic time of a single merging event can be written as $\tau_{h}\left(m_{h},a\right)={1/\left(\alpha_{0}f_{h}\right)}$.
Without loss of generality, the exponent $\lambda$ is a halo geometry parameter that represents the effect of the halo surface area on the merging frequency $f_{h}\left(m_{h},a\right)$. For two haloes of very different sizes (merging between a large halo and a single merger), it is estimated that $\lambda={2/3}$ with $m_{h}\propto r_{h}^{3}\propto A_{h}^{{3/2}}$, where $A_{h}$ is the surface area of the halo. For small haloes, merging is more likely between two haloes of comparable sizes, where $\lambda$ can deviate from 2/3 and approach 1.
Substitution of Eq. (145) for the event frequency into Eq. (144) leads to the mass flux,
@@
For self-similar gravitational clustering in the mass propagation range ($m_{h}<m_{h}^{*}$), the halo mass $m_{h}$ and the characteristic mass scale $m_{h}^{*}$ are the only two controlling variables. We can simply express the mass function as $f_{M}\left(m_{h},m_{h}^{*}\right)\sim\left(m_{h}\right)^{x}\left(m_{h}^{*}\right)^{-x-1}$. Now using dimensional analysis, the only possible form of the mass function $f_{M}$ that satisfies Eq. (146) is ($f_{M}$ should have a unit of 1/kg in SI units and $\varepsilon_{m}$ is a function of a alone and is independent of $m_{h}$),
@@
which is a constant invariant in time and dependent only on the mass resolution $m_{p}$. For $m_{p}=2.27\times 10^{11}{M_{\odot}/h}$ from Table 2 and $m_{h}^{*}\left(z=0\right)\approx 2\times 10^{13}{M_{\odot}/h}$, $\lambda_{0}$ is on the order of 0.13.
@@
that is also related to the Hubble constant $H\left(a\right)$ and the mass resolution $m_{p}$. The characteristic time scale $\tau_{h}^{*}=\tau_{h}\left(m_{h}^{*},a\right)$ associated with the characteristic mass $m_{h}^{*}$ is
@@
The fundamental frequency $f_{0}\left(a\right)$ is the frequency for the elementary merging of two single mergers and is expected to decrease with time. Without loss of generality, let us assume a power-law of $f_{0}\left(a\right)\propto a^{-\tau_{0}}$ that leads to $\varepsilon_{m}\left(a\right)\propto a^{-\tau_{0}}$ (Eq. (148)), $\left(m_{h}^{*}\right)^{1-\lambda}\propto a^{{3/2}-\tau_{0}}$ (Eq. (154)), and $M_{h}\left(a\right)\propto a^{{3/2}-\tau_{0}}$ (Eq. (150)). Once the statistically steady state is established for inverse mass cascade, the total mass of the halo sub-system increases as $M_{h}\left(a\right)\propto a^{{3/2}-\tau_{0}}$ regardless of the value of $\lambda$. Obviously,
@@
With the total mass in the halo sub-system $M_{h}\left(a\right)$ increasing with the scale factor a, it requires $0<\tau_{0}<{3/2}$. With $\tau_{0}>0$ and $a\to\infty$, the mass flux $\varepsilon_{m}\left(a\right)$ approaches zero when the whole system approaches the limiting thermodynamic equilibrium but cannot reach. In this regard, the inverse mass cascade with a constant rate is a key feature of the statistically (intermediate) steady state as the system evolves toward the limiting equilibrium.
The number of haloes $n_{h}\left(m_{h}\right)$ in the halo group with a given mass $m_{h}$ is
@@
, which does not vary with time once the statistically steady state is established. Substituting Eq. (157) into Eq. (155), we can express the fundamental frequency $f_{0}\left(a\right)$ as
@@
which is dependent on the mass resolution $m_{p}$ only (the mass of dark matter particle), with a smaller particle mass $m_{p}$ giving rise to a higher fundamental frequency $f_{0}\left(a\right)$ in the N-body simulation. For the mass resolution of $m_{p}=2.27\times 10^{11}{M_{\odot}/h}$ from Table 2 and $m_{h}^{*}\left(a=1\right)=2\times 10^{13}{M_{\odot}/h}$, $b_{0}\approx 2.2$ with $\tau_{0}=1$. In other words, the mass of dark matter particles can be determined if the fundamental frequency $f_{0}\left(a\right)$ can be precisely measured.
@@
which is the mass fraction of halo sub-system $M_{h}\left(a\right)$ and is not dependent on the mass resolution $m_{p}$ and the scale factor a. We estimate $c_{0}\approx 0.29$ with $\tau_{0}=1$, i.e. $M_{h}\approx 0.58Nm_{p}$ when a=1. This information is used to study the density distributions for particles in haloes and out-of-haloes, respectively (Xu:2022-Two-thirds-law-for-pairwise-ve).
We present simplified expressions for time and mass scales and mass flux function that can be fully described as functions of halo mass $m_{h}$, scale factor a, and mass resolution $m_{p}$ with four numerical constants $\beta_{0}$ (pre-factor for the halo mass function $f_{M}$), $c_{0}$ (the mass fraction of the halo sub-system), $\lambda$ (the halo geometry parameter) and $\tau_{0}$ (the exponent of the fundamental frequency $f_{0}$), and the characteristic mass $m_{h}^{*}\left(a=1\right)$ and $H_{0}$, with four
@@
Recall the time scales we introduced in Section C.2. The time scale $\tau_{M}\left(a\right)$ is
@@
The time scale $\tau_{h}$ for a single merging in a halo group of mass $m_{h}$ is,
@@
The time scale $\tau_{g}$ that takes to cascade the group mass $m_{g}$ for the halo group of mass $m_{h}$ is
@@
As expected, the mean waiting time (lifespan) $\tau_{g}$ of a given halo decreases with halo mass as $\tau_{g}\propto m_{h}^{-\lambda}$ and increases with a. Larger haloes have a shorter life. Extremely large haloes have very fast mass accretion and an infinitesimal lifespan.
@@
which increases with halo mass as $\tau_{f}\propto m_{h}^{1-\lambda}a^{\tau_{0}}$ such that Smaller haloes were formed earlier with smaller $\tau_{f}$. The relation between time scales $\tau_{f}\left(m_{h}^{*},a\right)=\beta_{0}\tau_{M}\left(a\right)$ can be easily obtained from Eqs. (166) and (169) that is consistent with our analysis in Section C.2 (Eq. (136)), i.e. $\beta_{0}M_{h}\left(a\right)=m_{g}^{*}n_{p}^{*}=n_{h}^{*}n_{p}^{*2}m_{p}$ and $z_{h}^{*}\beta_{0}=1$ (from Eq. (137)). Different time scales can be related to the mass flux as
@@
The corresponding time scales at the characteristic mass $m_{h}^{*}$ are
@@
The two time scales $\tau_{g}\propto a^{\tau_{0}}m_{h}^{-\lambda}$ (Eq. (168)) and $\tau_{f}\propto a^{\tau_{0}}m_{h}^{1-\lambda}$ (Eq. (169)), where larger haloes have a shorter lifespan but take more time to form. For $\lambda={2/3}$ and $\tau_{0}=1$, the lifespan $\tau_{g}^{*}$ of characteristic haloes is independent of time which is consistent with Eq. (32).
@@
which follows the same scaling as characteristic mass scale $m_{h}^{*}\left(a\right)$. For large haloes with $\tau_{0}=1$ and $\lambda={2/3}$, $m_{h}^{L}\left(a\right)\sim m_{h}^{*}\left(a\right)\sim a^{{3/2}}$. This scaling matches the mass growth of type II haloes, i.e., the dominant type for large haloes (see McBride:2009-Mass-accretion-rates-and-histo, Fig. 2), as shown in Fig. 23.
Figure 23: The halo mass (normalized by $10^{12}M_{\odot}$) accretion history for type II haloes, i.e. the dominant type for large haloes (see McBride:2009-Mass-accretion-rates-and-histo, Fig. 2), exhibits a power law scaling $\propto a^{{3/2}}$.
With the help of Eq. (133), we confirm that the time scale $\tau_{f}$ (Eq. (169)) to form the typical halo is on the same order of t (Eq. (133)),
@@
C.4 The probability distribution of waiting time